OpenStudy (anonymous):

im lazy will get.What is the equation of the line that contains the points (3, –6) and (–3, 0)?

OpenStudy (anonymous):

will get medal

OpenStudy (texaschic101):

I am too lazy to help

OpenStudy (anonymous):

Given points are: (3, –6) =(x1, y1) and (–3, 0)=(x2, y2) eQUATION OF LINE IN 2 POINT FORM IS GIVEN AS: y-y1= m(x-x1) m= (y1-y2)/(x1-x2) =(-6-0)/(3+3) =(-6)/(6) =-1 hence eq of line y-y1= m(x-x1) i.e. y-(-6)= -1(x-3) i.e. y +6= -x+3 x+y = -3 is the required eq of the line